# An Algorithmic Approach to Finding Degree-Doubling Nodes in Oriented Graphs

### v10

@@ 4 Initial Lemmas
What follows is a set of lemmas that seek to place a lower bound on the out-degree of a minimum out-degree node in a counterexample. Unless otherwise noted, we assume that $v_{0}$ is a node with the minimum out-degree in our oriented graph $G$. The work in this section agrees with what was already done by [10] and is not intended as original content. We provide it here to guide readers into our proof by contradiction methods. As we stated in the introduction, we begin with a single node, asking what requirements need to be in place for that node’s degree not to double. If we are able to successfully achieve a single node whose degree does not double, and list the accompanying requirements, we would like to continue to move to other nodes and the complete graph.
@@
Minimum Out-Degree $<$ 3
Suppose that the root minimum out-degree node $v_{0}$ in the oriented graph $G$ has an out-degree less than 3. Then $v_{0}$’s out-degree will at least double.
@@
Case 3: Assume that $d^{+}(v_{0})=2$. Let’s call the out-neighbors of $v_{0}$, $v_{1}$ and $v_{2}$, where $|N^{+}(v_{0})|=2$. We know that $d^{+}(v_{1})\geq 2$ and $d^{+}(v_{2})\geq 2$ since $v_{0}$ is a minimum out-degree node. Consider the induced subgraph $G[v_{1},v_{2}]$. At most one arc can exist in an oriented graph of size 2 because oriented graphs do not allow for symmetric arcs. So we know that there is at most one arc between $v_{1}$ and $v_{2}$ in this induced subgraph. This means that $v_{1}$ or $v_{2}$ will have at least two arcs outside of $G[v_{1},v_{2}]$. Neither $v_{1}$ or $v_{2}$ can relate to $v_{0}$ since $G$ is an oriented graph and $(v_{0},v_{1})\in G$ and $(v_{0},v_{2})\in G$. We can define the nodes $v_{3}$ and $v_{4}$ as the two out-neighbors of $v_{1}$ not in ${v_{0},v_{1},v_{2}}$. This means that $v_{0}$ will have $v_{3}$ and $v_{4}$ as second neighbors. Hence, $v_{0}$ will have at least two second out-neighbors—$v_{3}$ and $v_{4}$. Thus $|N^{++}(v_{0})|\geq 2$ and $|N^{++}(v_{0})|\geq|N^{+}(v_{0})|$, and $v_{0}$ is a degree-doubling node.
@@
Lemma 4.1 (Minimum Out-Degree $<$ 3) begins to tell us about the necessary conditions for degree doubling nodes in oriented graphs. What we see from this first lemma is that we need the minimum degree of our graph to be at least 3. In each of the cases of the lemma, the minimum out-degree node has very few out-neighbors, not enough to support the load balancing that we well see is necessary to prevent the degree doubling of nodes.
@@
Assume that $v_{1}$, which is a neighbor of $v_{0}$ has out-degree 0 in $G[v_{1},v_{2},v_{3}]$. This means that $(v_{1},v_{2})\notin G[v_{1},v_{2},v_{3}]$ and $(v_{1},v_{3})\notin G[v_{1},v_{2},v_{3}]$. Since $v_{0}$ is a minimum out-degree node with out-degree 3, we know that $v_{1}$ must have out-degree at least 3. So $v_{1}$ must connect to at least 3 out-neighbors outside of $\{v_{1},v_{2},v_{3}\}$. Since $(v_{0},v_{1})\in G$, we know that $(v_{1},v_{0})\notin G$ by the oriented property of $G$. This means that $v_{1}$ must connect to at least three other nodes $v_{4},v_{5},v_{6}$. So we have $|N^{+}_{G}(v_{1})|\geq 3$. The nodes $v_{4},v_{5},v_{6}$ are not in $G[v_{1},v_{2},v_{3}]z$ and are not equal to $v_{0}$, so they are outside $G[v_{0},v_{1},v_{2},v_{3}]$.
@@
Minimum Out-Degree 3 with Neighbors 1 in the Neighbor Induced Subgraph Assume that in an oriented graph $))G$, the minimal out-degree node $v_{0}$ has an out-degree of 3. Additionally, suppose that in the induced subgraph of $v_{0}$’s out-neighbors, every neighbor has an out-degree of 1. Then the out-degree of $v_{0}$ or one of its neighbors will be a degree-doubling node.
@@
This reduces the situation to Case 3 in Lemma 4.1 (Minimum Out-Degree $<$ 3), where $v_{1}$ acts similarly to how $v_{0}$ did in that case. There is at most one arc between $v_{4}$ and $v_{5}$. We’ll say $v_{4}$ will have three additional out-neighbors, $v_{6},v_{7}$, and $v_{8}$, with degree zero in $G[v_{4},v_{5}]$. Then $|N^{+}(v_{4})|=3$. These three neighbors of $v_{4}$ are second out-neighbors of $v_{1}$, giving us $|N^{++}(v_{1})|=3$. Thus we have the following equality, $|N^{++}(v_{1})|=|N^{+}(v_{1})|=3$, making $v_{1}$ a degree-doubling node.
Hence, $v_{1}$ must be a degree-doubling node, contradicting the assumption that such a node could exist in a minimal counterexample.
Lemma 4.3 (Minimum Out-Degree 3 with Neighbors 1) hints at a strategy of examining the out-neighborhoods of neighbors, which is a question we shall revisit throughout this paper. It must prevent a node’s out-degree from doubling for it to function. In this instance, we observe that the degree of the starting node in the example $v_{0}$ is not doubled. We are compelled to look for the next node instead. In this case, the degree of the neighbor of $v_{0}$, $v_{1}$, doubles.
We did not require the right starting node. Therefore, this is a crucial distinction. Not because of its possible proximity to the degree-doubling node. Instead, we chose a minimum out-degree node due to its ability to split the graph’s nodes. We can see an illustration of Lemma 4.3 in Example 4.2.
@@
5 The GLOVER (Graph Level Order)
In the world of computer science, data structures are a fundamental tool for organizing information. Efficient data structures have helped not only provide much of how the real world works today. Not only that, but many of the major open problems were helped with data structures. The development of balanced trees (AVL trees, Red-Black trees, and B-trees) helped with efficient searching and information retrieval. Graph data structures have been crucial for algorithms like max flow/min cut problems and network optimization problems. Stacks, queues and parse trees help with compiler design and programming languages.
We have begun a bottom up approach to the Seymour Second Neighborhood Conjecture, starting with a minimum degree node. What we are going to do is continue this approach, culminating in a data structure that will order the nodes by distance from that minimum degree node. Conceptually, this is how Example 4.2 was drawn, but we will make this a part of our formal notation. By doing this, we will be able to take advantage of other properties of this data structure and its organization of the nodes within it.
The Example 4.2 is not the first time we have seen the minimum out-degree node not having its degree double. Fisher also provided an example in [8] where the minimum out-degree node does not have its degree double. In Example 4.2, this node happens to be an out-neighbor of the minimum out-degree node. There is no lasting guarantee that this will always be the case. There is no reason why future examples would have to stop in first out-neighbors of minimum out-degree nodes. This process could continue through a sequence of out-neighbors of the minimum out-degree node. No proof has been given that this process has to stop anywhere just yet.
Up until now, we have been using the terminology of induced subgraphs to represent adjacent out-neighborhoods. We have just begun to speak of why this might not be sufficient for what we have planned. In Example 4.2, it was not difficult to speak of the single induced subgraph and the next induced subgraph using $G[v_{1},v_{2},v_{3}]$ and $G-[v_{1},v_{2},v_{3}]$, respectively. If this degree-doubling node had been a further distance from $v_{0}$, expressing it through such an induced subgraph representation would not have been so easy. As the distance between the degree-doubling node and the minimum out-degree node $v_{0}$ is increased, the number of induced subgraphs begins to become longer, and this notation becomes challenging. Instead, we plan to introduce a new terminology that covers our thought process.
@@ 5.1 Definitions
In this section, we develop a new data structure for oriented graphs, Graph Level Order (GLOVER). This structure is specifically designed for the SSNC and oriented graphs but will have applications beyond. In order to define this data structure, we will need to formally refer to the partitions we discussed earlier as rooted neighborhoods. Additionally, we shall discuss interior, exterior, and back arcs and show the influence they have on our data structure. Finally, we will speak about how this data structure can use the path-finding technique to help solve the SSNC.
@@
Given an oriented graph $G$ and a minimum out-degree node $v_{0}$, a rooted neighborhood $R_{i}$ of distance $i$ is the subgraph of $G$ induced by the set of all nodes at distance exactly $i$ from $v_{0}$. Formally, let $V_{i}(v_{0})=\{v\in V(G)\mid\text{dist}(v_{0},v)=i\}$. Then,
$R_{i}=G[V_{i}(v_{0})].$
Example 5.1.
Figure 6: Illustration of rooted neighborhood. Rooted neighborhoods group nodes based on their distance from the minimum out-degree node $v_{0}$, with no assumptions on the sizes of these rooted neighborhoods.
Similar constructions have appeared in prior work. For instance, Daamouch employed induced subgraphs in his analysis of $m$-free graphs [1]. However, the focus here is more sharply placed on the hierarchical structure of rooted neighborhoods (see Example 5.1), which plays a central role in our proofs. This is also closely related to the notion of the $i$-th rooted neighborhood introduced independently by Botler, Moura, and Naia [3], who defined neighborhoods in terms of shortest paths. Their work focused on the global behavior of (pseudo)random graph orientations. While their results are powerful and address large-scale properties, our aim is to explore the fine-grained, local behavior of these rooted neighborhoods. We propose that understanding these local node interactions will give rise to clearer global insights of the conjecture.
@@
These rooted neighborhoods are anchored at a minimum out-degree node, and their importance is revealed by the element $i$, which identifies the distance between that rooted neighborhood and this minimum out-degree node. This allows us to not only partition the nodes, which we have already done by the rooted neighborhoods, but now we can keep track of entire paths from $v_{0}$ to every node in every $u_{i}\in R_{i}$. These nodes within the rooted neighborhoods will have hierarchical relationships. That is, relationships with nodes in the rooted neighborhoods closer to $v_{0}$, relationships with nodes in their own rooted neighborhood, and relationships with nodes in rooted neighborhoods further from $v_{0}$. These relationships are the backbone of the GLOVER data structure. Within this section, we will show that this is not just a labeling. There are important properties that are revealed within these rooted neighborhoods being at distance $i$ from the minimum out-degree node.
At the heart of these rooted neighborhoods are parent-child relationships. A node $u_{i}$ at level $R_{i}$ has out-neighbors at level $R_{i+1}$. These out-neighbors at the next level $v_{1},v_{2},\ldots,v_{n}$, which are the children of $u_{i}$, have a sibling relationship to each other. In this proof by contradiction, the goal of the parent at each level $R_{i}$ is to prevent their degree from doubling. Demonstrating how their children, through a process of interlocking, help share that load will show this is impossible. The next section formalizes this, detailing the graph-theoretic specifics of the interlocking and how it prevents degree doubling. First, however, the meaning of ”parent,” ”child,” ”sibling,” and ”the load” in this context will be clarified.
@@
Transitive Triangle Let $(x,y),(x,u),(y,u)\in G$. Then $x$, $y$, and $u$ form a transitive triangle, where $x$ is a common predecessor of both $y$ and $u$, and $y$ also connects to $u$.
@@
Parent and Child Let $u_{i}$ be a node in the rooted neighborhood $R_{i}$ for some $i\geq 0$. A child of $u_{i}$ is a node $v_{i+1}\in R_{i+1}$ such that $(u_{i},v_{i+1})\in G$ (we also say that $u_{i}$ is the parent of $v_{i+1}$).
@@
$deg_{\text{int}}(u_{i},v_{1}):=|\text{int}(u_{i},v_{1})|$
@@
With the concept of transitive triangles now defined in the context of parent, child, and child-neighbor relationships, we can begin to examine the role of children in supporting their parents. To approach the problem from a proof by contradiction angle, we consider a start node that we would like to avoid having its degree double. Rather than describing this in strict mathematical terms, we can frame it metaphorically: the node (or parent) does not want to carry a heavy burden. We think of this as something that could potentially overwhelm it.
@@
Because these second neighbors are akin to the counterparts to the first neighbors, they should have relationships to those same parent $u$ and child $v$ nodes in the transitive triangle, but from a different angle. They are second neighbors of the parent node $u$, so the conjecture itself will bound the size of these elements. Likewise, these exterior neighbors will be related to the child node $v$ because there must be a path from $u\rightarrow v\rightarrow w$ for any exterior neighbor $w$.
While the interior neighbors portrays a concept similar to transitive triangles, the notion of exterior neighbors is entirely new. The core of the idea is that once we have these rooted neighborhoods, a node’s degree can be split into interior and exterior degrees. Within those two degrees, we expect two different types of behaviors. One type of behavior, interior degree, we expect to be able to show is easy, predictable, and tractable behavior. This is because these interior neighbors, relative to the parent nodes, will cause cycles among the interior neighbors to prevent that parent from becoming a degree doubling node. Exterior degree, on the other hand, will be a more strenuous effort. To understand how exterior neighbors impact the conjecture, we need to see how they interact with the GLOVER data structure and the Decreasing Neighborhood Sequence Property (DNSP).
@@
The concept of interior neighbors, as we define it, represents a specific and crucial type of transitive triangle that arises within our GLOVER structure. While Brantner et al. [2] were instrumental in highlighting the general significance of transitive triangles—which can also arise from back arcs and present hindrances to the conjecture and exterior neighbors, which do not present a problem to the conjecture—and introduced the related idea of shared first out-neighbors, our formulation introduces a critical distinction. Our definition of interior neighbors specifically focuses on the directed arcs between the children of a common parent within the GLOVER structure and how they relate to that parent. Furthermore, Brantner et al.’s work did not extend to a parallel formulation for exterior neighbors.
We postulate that, much like the necessity of studying both primal and dual programs to achieve a comprehensive understanding in linear programming, a complete resolution of the Seymour Second Neighborhood Conjecture requires the dual analysis of both interior and exterior neighbors. As we will demonstrate throughout this paper, the properties of these interior and exterior neighbors—which are fundamentally defined by their relationship to a common parent within the GLOVER structure—will provide us with crucial bounds on the actions and limitations of that parent node and its implications for the conjecture.
Importantly, it is worth noting a further subtlety regarding back arcs. The data structure itself does not handle these arcs well. Instead, our traversal algorithm will be able to treat these arcs as exterior arcs, and quickly locate a degree doubling node.
@@
For any $v\in R_{i+1}$ with parent $u\in R_{i}$, the out-neighbors of $v$ can be partitioned as:
$N^{+}(v)=\text{int}(u,v)\cup\text{ext}(u,v)\cup back(v)$
@@
Let $v\in R_{i+1}$ be a node with parent $u\in R_{i}$. We define the following three sets for any out-neighbor $w\in N^{+}(v)$:
@@
  • $\text{int}(u,v)=\{w\in N^{+}(v)\mid\text{dist}(v_{0},w)=\text{dist}(v_{0},v)\land(u,w)\in G\}$
  • $\text{ext}(u,v)=\{w\in N^{+}(v)\mid(\text{dist}(v_{0},w)=\text{dist}(v_{0},v)\land(u,w)\notin G)\lor(\text{dist}(v_{0},w)>\text{dist}(v_{0},v))\}$
@@
Case 1: $\text{dist}(v_{0},w)<\text{dist}(v_{0},v)$ By definition, $w\in\text{back}(v)$. Such an arc $(v,w)$ is a back arc in the rooted neighborhood structure. Since its distance from $v_{0}$ is less than $\text{dist}(v_{0},v)$, it cannot satisfy the conditions for $\text{int}(u,v)$ or $\text{ext}(u,v)$.
@@
  • If $(u,w)\in G$, then $w\in\text{int}(u,v)$. This means $w$ is an out-neighbor of $v$ within the same layer, and also a direct out-neighbor of $v$’s parent $u$.
  • If $(u,w)\notin G$, then $w\in\text{ext}(u,v)$. This means $w$ is an out-neighbor of $v$ within the same layer, but not a direct out-neighbor of $u$.

### v12

@@ 4 Initial Lemmas
What follows is a set of lemmas that seek to place a lower bound on the out-degree of a minimum out-degree node in a counterexample. Unless otherwise noted, we assume that $v_{0}$ is a node with the minimum out-degree in our oriented graph $G$. The work in this section agrees with what was already done by [20] and is not intended as original content. We provide it here to guide readers into our proof by contradiction methods. As we stated in the introduction, we begin with a single node, asking what requirements need to be in place for that node’s degree not to double. If we are able to successfully achieve a single node whose degree does not double, and list the accompanying requirements, we would like to continue to move to other nodes in the graph.
@@
Minimum Out-Degree $<$ 3 Suppose that the minimum out-degree node $v_{0}$ in the oriented graph $G$ has an out-degree less than 3. Then $v_{0}$’s out-degree will at least double.
@@
Case 3: Assume that $d^{+}(v_{0})=2$. Let’s call the out-neighbors of $v_{0}$, $v_{1}$ and $v_{2}$, where $|N^{+}(v_{0})|=2$. We know that $d^{+}(v_{1})\geq 2$ and $d^{+}(v_{2})\geq 2$ since $v_{0}$ is a minimum out-degree node. Consider the induced subgraph $G[v_{1},v_{2}]$. At most one arc can exist in an oriented graph of size 2 because oriented graphs do not allow for symmetric arcs. So we know that there is at most one arc between $v_{1}$ and $v_{2}$ in this induced subgraph. This means that $v_{1}$ or $v_{2}$ will have at least two arcs outside of $G[v_{1},v_{2}]$. Neither $v_{1}$ or $v_{2}$ can relate to $v_{0}$ since $G$ is an oriented graph and $v_{0}\to v_{1}\in G$ and $v_{0}\to v_{2}\in G$. We can define the nodes $v_{3}$ and $v_{4}$ as the two out-neighbors of $v_{1}$ not in ${v_{0},v_{1},v_{2}}$. This means that $v_{0}$ will have $v_{3}$ and $v_{4}$ as second neighbors. Hence, $v_{0}$ will have at least two second out-neighbors—$v_{3}$ and $v_{4}$. Thus $|N^{++}(v_{0})|\geq 2$ and $|N^{++}(v_{0})|\geq|N^{+}(v_{0})|$, and $v_{0}$ is a degree-doubling node.
@@
Lemma 4.1 (Minimum Out-Degree $<$ 3) begins to tell us about the necessary conditions for degree doubling nodes in oriented graphs. What we see from this first lemma is that we need the minimum degree of our graph to be at least 3. In each of the cases of the lemma, the minimum out-degree node has very few out-neighbors, not enough to support the load balancing that we will see is necessary to prevent the degree doubling of nodes.
@@
Assume that $v_{1}$, which is a neighbor of $v_{0}$ has out-degree 0 in $G[v_{1},v_{2},v_{3}]$. This means that $v_{1}\to v_{2}\notin G[v_{1},v_{2},v_{3}]$ and $v_{1}\to v_{3}\notin G[v_{1},v_{2},v_{3}]$. Since $v_{0}$ is a minimum out-degree node with out-degree 3, we know that $v_{1}$ must have out-degree at least 3. So $v_{1}$ must connect to at least 3 out-neighbors outside of $\{v_{1},v_{2},v_{3}\}$. Since $v_{0}\to v_{1}\in G$, we know that $v_{1}\to v_{0}\notin G$ by the oriented property of $G$. This means that $v_{1}$ must connect to at least three other nodes $v_{4},v_{5},v_{6}$. So we have $|N^{+}_{G}(v_{1})|\geq 3$. The nodes $v_{4},v_{5},v_{6}$ are not in $G[v_{1},v_{2},v_{3}]$ and are not equal to $v_{0}$, so they are outside $G[v_{0},v_{1},v_{2},v_{3}]$.
@@
Minimum Out-Degree 3 with Neighbors 1 in the Neighbor Induced Subgraph Assume that in an oriented graph $G$, the minimal out-degree node $v_{0}$ has an out-degree of 3. Additionally, suppose that in the induced subgraph of $v_{0}$’s out-neighbors, every neighbor has an out-degree of 1. Then the out-degree of $v_{0}$ or one of its neighbors will be a degree-doubling node.
@@
This reduces the situation to Case 3 in Lemma 4.1 (Minimum Out-Degree $<$ 3), where $v_{1}$ acts similarly to how $v_{0}$ did in that case. There is at most one arc between $v_{4}$ and $v_{5}$. We will say that $v_{4}$ has three additional out-neighbors, $v_{6},v_{7}$, and $v_{8}$, with $v_{4}$ having degree zero in $G[v_{4},v_{5}]$. Then $|N^{+}(v_{4})|=3$. These three neighbors of $v_{4}$ are second out-neighbors of $v_{1}$, giving us $|N^{++}(v_{1})|=3$. Thus we have the following equality, $|N^{++}(v_{1})|=|N^{+}(v_{1})|=3$, making $v_{1}$ a degree-doubling node.
Hence, $v_{1}$ must be a degree-doubling node, contradicting the assumption that such a node could exist in a minimal counterexample. ∎
Lemma 4.3 (Minimum Out-Degree 3 with Neighbors 1 in the Neighbor Induced Subgraph) hints at a strategy of examining the out-neighborhoods of neighbors, which is a question we shall revisit throughout this paper. It must prevent a node’s out-degree from doubling for it to function. In this instance, we observe that the degree of the starting node in the example $v_{0}$ is not doubled. We are compelled to look for the next node instead. In this case, the degree of the neighbor of $v_{0}$, $v_{1}$, doubles.
We did not require the right starting node. Therefore, this is a crucial distinction. Not because of its possible proximity to the degree-doubling node. Instead, we chose a minimum out-degree node due to its ability to split the graph’s nodes. We can see an illustration of Lemma 4.3 (Minimum Out-Degree 3 with Neighbors 1 in the Neighbor Induced Subgraph) in Example 4.2.
@@
5 Graph Level Order
In the world of computer science, data structures are a fundamental tool for organizing information. Efficient data structures have helped not only provide much of how the real world works today. Not only that, but many of the major open problems were helped with data structures. The development of balanced trees (AVL trees, Red-Black trees, and B-trees) helped with efficient searching and information retrieval. Graph data structures have been crucial for algorithms like max flow/min cut problems and network optimization problems. Stacks, queues and parse trees helped with compiler design and programming languages. Data structures have also been front and center in the world of open problems.
We have begun a bottom up approach to the SSNC, starting with a minimum degree node. What we are going to do is continue this approach, culminating in a data structure that will order the nodes by distance from that minimum degree node. Conceptually, this is how Example 4.2 was drawn, but we will make this a part of our formal notation. By doing this, we will be able to take advantage of other properties of this data structure and its organization of the nodes within it.
The Example 4.2 is not the first time we have seen the minimum out-degree node not having its degree double. Fisher also provided an example in [11] where the minimum out-degree node does not have its degree double. In Example 4.2, this node happens to be a first out-neighbor of the minimum out-degree node. There is no lasting guarantee that this will always be the case. There is no reason why future examples would have to stop in first out-neighbors of minimum out-degree nodes. This process could continue through a sequence of out-neighbors of the minimum out-degree node. No proof has been given that this process has to stop anywhere just yet.
Up until now, we have been using the terminology of induced subgraphs to represent adjacent out-neighborhoods. We have just begun to speak of why this might not be sufficient for what we have planned. In Example 4.2, it was not difficult to speak of the single induced subgraph and the next induced subgraph using $G[v_{1},v_{2},v_{3}]$ and $G-[v_{1},v_{2},v_{3}]$, respectively. If this degree-doubling node had been a further distance from $v_{0}$, expressing it through such an induced subgraph representation would not have been so easy. As the distance between the degree-doubling node and the minimum out-degree node $v_{0}$ is increased, the number of induced subgraphs begins to become longer, and this notation becomes challenging. What might begin to happen is that notation becomes confusing or intent begins to get in the way over innovation, so time spent discussing ideas for a proof is instead spent trying to understand parentheses to ensure that we are all in the correct ’induced subgroup’. Instead, we plan to introduce a new terminology that covers our thought process, and ’keep the main thing the main thing’.
@@ 5.1 Definitions
In this section, we develop a new data structure for oriented graphs, Graph Level Order. This structure is specifically designed for the SSNC and oriented graphs but will have applications beyond. In order to define this data structure, we will need to formally refer to the partitions we discussed earlier as rooted neighborhoods. Additionally, we shall formally define interior, exterior, and back arcs and show the influence they have on our data structure. Finally, we will speak about how this data structure can use the path-finding technique to help solve the SSNC.
@@
Given an oriented graph $G$ and a minimum out-degree node $v_{0}$, a rooted neighborhood $R_{i}$ of distance $i$ is the subgraph of $G$ induced by the set of all nodes at distance exactly $i$ from $v_{0}$. Formally, let
$N_{i}^{+}(v_{0})=\{v\in G\mid\text{dist}(v_{0},v)=i\}$
Then, $R_{i}=G[N_{i}^{+}(v0)]$.
Figure 6: Illustration of a rooted neighborhood. Rooted neighborhoods group nodes based on their distance from the minimum out-degree node $v_{0}$, with no assumptions on the sizes of these rooted neighborhoods.
Similar constructions have appeared in prior work. For instance, Daamouch employed induced subgraphs in his analysis of $m$-free graphs [6]. However, the focus here is more sharply placed on the hierarchical structure of rooted neighborhoods (see Example 6), which plays a central role in our proofs. This is also closely related to the notion of the $i$-th rooted neighborhood introduced independently by Botler et al. [1], who defined neighborhoods in terms of shortest paths. Their work focused on the global behavior of (pseudo)random graph orientations. While their results are powerful and address large-scale properties, our aim is to explore the fine-grained, local behavior of these rooted neighborhoods. We propose that understanding these local node interactions will give rise to clearer global insights of the conjecture.
@@
These rooted neighborhoods are anchored at a minimum out-degree node, and their importance is revealed by the element $i$, which identifies the distance between that rooted neighborhood and this minimum out-degree node. This allows us to not only partition the nodes, which we have already done by the rooted neighborhoods, but now we can keep track of entire paths from $v_{0}$ to every node in every $u_{i}\in R_{i}$. These nodes within the rooted neighborhoods will have hierarchical relationships. That is, relationships with nodes in the rooted neighborhoods closer to $v_{0}$, relationships with nodes in their own rooted neighborhood, and relationships with nodes in rooted neighborhoods further from $v_{0}$. These relationships are the backbone of the Graph Level Order. Within this section, we will show that this is not just a labeling. There are important properties that are revealed within these rooted neighborhoods being at distance $i$ from the minimum out-degree node.
At the heart of these rooted neighborhoods are parent-child relationships. A node $u_{i}$ at level $R_{i}$ has out-neighbors at level $R_{i+1}$. These out-neighbors at the next level $v_{1},v_{2},\ldots,v_{n}$, which are the children of $u_{i}$, have a sibling relationship to each other. In this proof by contradiction, the goal of the parent at each level $G[N^{i+}]$ is to prevent their degree from doubling. Demonstrating how their children, through a process of interlocking, help share that load will show this is impossible. The next section formalizes this, detailing the graph-theoretic specifics of the interlocking and how it prevents degree doubling. First, however, the meaning of ”parent,” ”child,” ”sibling,” and ”the load” in this context will be clarified.
@@
Transitive Triangle Let $x\to y,x\to u,y\to u$. Then $x$, $y$, and $u$ form a transitive triangle.
@@
Parent and Child Let $u_{i}$ be a node in the rooted neighborhood $R_{i}$ for some $i\geq 0$. A child of $u_{i}$ is a node $v_{i+1}\in R_{i+1}]$ such that $u_{i}\to v_{i+1}$ (we also say that $u_{i}$ is the parent of $v_{i+1}$).
@@
Interior Neighbor and Interior Degree Let $u_{i}\in R_{i}$ be a parent node with children $v_{1},v_{2}\in R_{i+1}$. We define the interior neighbors of $v_{1}$ with respect to $u_{i}$ as those nodes $z\in G$ such that both $u_{i}\to z$ and $v_{1}\to z$. That is, nodes that are common out-neighbors of both $u_{i}$ and $v_{1}$, forming transitive triangles.
@@
The interior degree of $v_{1}$ with respect to $u_{i}$ is defined as
$|\text{int}(u_{i},v_{1})|$
@@
With the concept of transitive triangles now defined in the context of parent, child, and child-neighbor relationships, we can begin to examine the role of children in supporting their parents. To approach the problem from a proof by contradiction angle, we consider a start node that we would like to avoid having its degree double. Rather than describing this in strict mathematical terms, we can frame it metaphorically: the node (or parent) does not want to carry a heavy burden. We think of this as something that could potentially overwhelm it, like becoming a degree-doubling node.
@@
Because these second neighbors are akin to the counterparts to the first neighbors, they should have relationships to those same parent $u$ and child $v$ nodes in the transitive triangle, but from a different angle. They are second neighbors of the parent node $u$, so the conjecture itself will bound the size of these elements. Likewise, these exterior neighbors will be related to the child node $v$ because there must be a path from $u\to v\to w$ for any exterior neighbor $w$.
While the interior neighbors portrays a concept similar to transitive triangles, the notion of exterior neighbors is entirely new. The core of the idea is that once we have these rooted neighborhoods, a node’s degree can be split into interior and exterior degrees. Within those two degrees, we expect two different types of behaviors. One type of behavior, interior degree, we expect to be able to show is easy, predictable, and tractable behavior. This is because these interior neighbors, relative to the parent nodes, will cause cycles among the interior neighbors to prevent that parent from becoming a degree doubling node. Exterior degree, on the other hand, will be a more strenuous effort. To understand how exterior neighbors impact the conjecture, we need to see how they interact with the Graph Level Order and the Decreasing Neighborhood Sequence Property (DNSP).
@@
Let $u_{i}\in R_{i}$ be a parent of a node $v_{i+1}\in R_{i+1}$. The exterior neighbors of $v_{i+1}$ with respect to $u_{i}$ are nodes $z$ such that $z$ is a second out-neighbor of $u_{i}$ and a first out-neighbor of $v_{i+1}$, i.e., $z\in N^{++}(u_{i})\cap N^{+}(v_{i+1})$. This implies that there exists a path $u_{i}\to w\to z$, and an arc $v_{i+1}\to z$ exists, but $u_{i}\to z\notin G$. Unlike the interior neighbors, exterior neighbors are neighbors of the child that are not shared by the parent.
The exterior degree of $v_{i+1}$ with respect to $u_{i}$ is defined as
$|\text{ext}(u_{i},v_{i+1})|.$
@@
The concept of interior neighbors, as we define it, represents a specific and crucial type of transitive triangle that arises within our Graph Level Order. While Brantner et al. [2] were instrumental in highlighting the general significance of transitive triangles—which can also arise from back arcs and present hindrances to the conjecture and exterior neighbors, which do not present a problem to the conjecture—and introduced the related idea of shared first out-neighbors, our formulation introduces a critical distinction. Our definition of interior neighbors specifically focuses on the directed arcs between the children of a common parent within the Graph Level Order and how they relate to that parent. Furthermore, Brantner et al.’s work did not extend to a parallel formulation for exterior neighbors.
We postulate that, much like the necessity of studying both primal and dual programs to achieve a comprehensive understanding in linear programming [24], a complete resolution of the SSNC requires the dual analysis of both interior and exterior neighbors. As we will demonstrate throughout this paper, the properties of these interior and exterior neighbors—which are fundamentally defined by their relationship to a common parent within the Graph Level Order—will provide us with crucial bounds on the actions and limitations of that parent node and its implications for the conjecture.
Exterior neighbors do relate to what Brantner et at. [2] referred to as Seymour diamonds. These are subgraphs where a node $x$ connects to a node $z$ through two different paths $y_{1}$ and $y_{2}$, $x\to y_{1}\to z$ and $x\to y_{2}\to z$. Seymour diamonds presented a problem for the SSNC because the dual paths to the node $z$ meant that it could be improperly counted as two second neighbors of the node $x$.
This concept is related to exterior neighbors because in the Graph Level Order, the node $x$ would be a parent node, with the children $y_{1}$ and $y_{2}$ who are siblings. These children can then serve as an exterior neighbor to $z$, (i.e $\text{ext}(x,y_{1})=z$ and $\text{ext}(x,y_{2})=z$. This lets us know that $z$ is a second neighbor of $x$, but with trackers through the nodes $y_{1}$ and $y_{2}$, which resolves the issue of over-counting $z$.
Figure 7: Illustration of interior and exterior neighbors. Node 00 is the parent node. The nodes $1\to 2\to 3\to 1$ forming a cycle are the first neighbors (or children) of node 00. The interior neighbors are always distance one from the parent node. The nodes $\{4,5\}$ are the exterior neighbors from any arc that parent node 00 and a node in $R_{1}$. The exterior neighbors are always second neighbors of the parent node.
@@
Let $v_{0}$ be a minimum out-degree node. Suppose that $x$ is a node in the rooted neighborhood $R_{i}$. A back arc is defined as an arc $x\to y$ such that $y\in N^{+}(x)$ and $y\in R_{j}$, where $j<i$.
Figure 8: Illustration of a node in rooted neighborhood $R_{3}$ with a back arc to the rooted neighborhood $R_{1}$.
@@
Partition of Node’s Degree For any $v\in R_{i+1}$ with parent $u\in R_{i}$, the out-neighbors of $v$ can be partitioned as:
$N^{+}(v)=\text{int}(u,v)\cup\text{ext}(u,v)\cup\text{back}(v)$
@@
Let $v\in R_{i+1}$ be a node with parent $u\in R_{i}$. We have defined the following three sets for any out-neighbor $w\in N^{+}(v)$:
@@
  • $\text{int}(u,v)=\{w\in N^{+}(v)\mid\text{dist}(v_{0},w)=\text{dist}(v_{0},v)\land u\to w\in G\}$
  • $\text{ext}(u,v)=\{w\in N^{+}(v)\mid(\text{dist}(v_{0},w)=\text{dist}(v_{0},v)\land u\to w\notin G)\lor(\text{dist}(v_{0},w)>\text{dist}(v_{0},v))\}$
@@
Case 1: $\text{dist}(v_{0},w)<\text{dist}(v_{0},v)$ By definition, $w\in\text{back}(v)$. Such an arc $v\to w$ is a back arc in the rooted neighborhood structure. Since its distance from $v_{0}$ is less than $\text{dist}(v_{0},v)$, it cannot satisfy the conditions for $\text{int}(u,v)$ or $\text{ext}(u,v)$.
@@
  • If $u\to w\in G$. This means $w$ is an out-neighbor of $v$ within the same layer, and also a direct out-neighbor of $v$’s parent $u$. So $w\in\text{int}(u,v)$
  • If $u\to w\notin G$. This means $w$ is an out-neighbor of $v$ within the same layer, but not a direct out-neighbor of $u$. So $w\in\text{ext}(u,v)$